OpenStudy (anonymous):

A 87 arrow is fired from a bow whose string exerts an average force of 105 on the arrow over a distance of 75 . What is the speed of the arrow as it leaves the bow? What is the speed of the arrow as it leaves the bow? What is the speed of the arrow as it leaves the bow?

OpenStudy (anonymous):

87 wat? 87N, 87cm?

OpenStudy (anonymous):

It is 87g and the distance is 75cm and the force is 105N

OpenStudy (anonymous):

work = delta energy = 1/2mV^2 F*d = 1/2mV^2 sqrt(2F*d/m) = V

OpenStudy (anonymous):

mass in kg, distance in meters, Force in Newtons... speed will be in meters/s

OpenStudy (anonymous):

Hey I solved it for V and got 1.345m/s but it is not right. Did I solve it wrong?

OpenStudy (anonymous):

m= 87g * 1kg/1000g d = 75cm * 1m/100cm

OpenStudy (anonymous):

Oh ok I forget to convert grams. That should wrk

OpenStudy (anonymous):

Thx